OpenStudy (balala):

How do I solve this equation? 66 = 14/13x – 6

OpenStudy (skullpatrol):

$$\Huge 66=\frac{14}{13}x-6$$

OpenStudy (balala):

That is what I'm trying to solve

OpenStudy (skullpatrol):

When you say the word "solve" you must also say WHAT you what to solve for?

OpenStudy (balala):

I think I'm supposed to solve for x.

OpenStudy (skullpatrol):

Yes, that means you are trying to transform the equation into a simpler equation whose solution or solutions can easily be seen.

OpenStudy (balala):

Ok, am I supposed to make the improper fraction into a mixed number? I don't know what to do first..

OpenStudy (skullpatrol):

We use the idea of balance to solve equations, whatever you do to one side, you must also do the same thing to the other to keep it balanced.

OpenStudy (skullpatrol):

I would start by adding 6 to both sides.

OpenStudy (balala):

Ok, so it would be something like this? 72 = 14/13x Did I do it right?

OpenStudy (skullpatrol):

Yes, now multiply both sides by 13/14.

OpenStudy (balala):

So it would be 1008/13 = x

OpenStudy (skullpatrol):

$$\Huge 72*\frac{13}{14}=x$$

OpenStudy (balala):

Actually I think it is the improper fraction 14 over 13

OpenStudy (skullpatrol):

what is the improper fraction 14 over 13?

OpenStudy (balala):

You wrote in the equation above 72 * 13 over 14 = x I think it is supposed to be 72 * 14 over 13 = x

OpenStudy (skullpatrol):

What did you multiply both sides of the equation with?

OpenStudy (balala):

I'm confused now.. it looks like I multiplied it by 13 over 14.

OpenStudy (skullpatrol):

$$\Huge 72 = \frac{14}{13}x$$

OpenStudy (skullpatrol):

$$\Huge\frac{13}{14}* 72 = \frac{13}{14}*\frac{14}{13}x$$

OpenStudy (skullpatrol):

$$\Huge\frac{13}{14}* 72 = 1*x$$

OpenStudy (skullpatrol):

$$\Huge\frac{a}{b}* \frac{b}{a}=1$$

OpenStudy (balala):

So then it would be: 13/14 * 72/1 = x 936 = x

OpenStudy (balala):

Oh wait it would be 935 over 14 = x Is that right?

OpenStudy (skullpatrol):